### v1

# Bi-Lipschitz Smoothing under Ricci and Injectivity Bounds

Abstract
We prove that a complete Riemannian manifold with a positive uniform lower bound on injectivity radius and a positive uniform lower bound on Ricci curvature admits an $L^{\infty}$-close (bi-Lipschitz) smooth metric with two-sided Ricci curvature bounds and a uniform positive lower bound on injectivity radius. This answers Question 2 in the Morgan–Pansu list of open problems from the conference Modern Trends in Differential Geometry (São Paulo, 2018), proposed by L. Bandara. In the proof, we rely on controlled smoothing [2] with Croke’s universal local volume lower bound [3] and the Cheeger–Gromov–Taylor injectivity radius estimate [4].
1 Introduction
In this note, we work with smooth manifolds without boundary. Let $(M^{n},g)$ be a complete Riemannian manifold with $\operatorname{inj}(M,g)\geq\ell>0$ and $\operatorname{Ric}_{g}\geq k\,g$ for some $k>0$. Our goal is to form a bi-Lipschitz nearby smooth metric $h$ with a two-sided Ricci bound and a uniform positive injectivity radius bound. More precisely, we solve the following problem.
Question 2 (due to Bandara [1]). For every $\ell,k>0$, there exist $C,L,K>0$ with the following effect. Let $(M,g)$ be a complete Riemannian manifold with injectivity radius $inj(M,g)\geq\ell$ and Ricci curvature $Ric(g)\geq k$. Then there exists a metric $h$ on $M$ with $inj(M,h)\geq L$ and $|Ric(h)|\leq K$ such that
$\frac{1}{C}g\leq h\leq Cg.$
The proof below is built on known results, namely, controlled smoothing [2], Croke’s universal local volume lower bound [3], and the Cheeger–Gromov–Taylor injectivity radius estimate [4].
2 Main result
Theorem 1 (Smoothing under $\operatorname{inj}$ and $\operatorname{Ric}$ lower bounds).
We restate Question 2 in more detail. Let us fix an integer $n\geq 2$. For every $\ell>0$ and $k>0$, there exist constants $C,L,K>0$ (that depend only on $n,\ell,k$) with the following property.
Let $(M^{n},g)$ be a complete Riemannian manifold that satisfies:
$\operatorname{inj}(M,g)\geq\ell,\qquad\operatorname{Ric}_{g}\geq k\,g\quad\text{(as quadratic forms).}$
There exists a smooth complete Riemannian metric $h$ on $M$ with:
$\frac{1}{C}\,g\leq h\leq C\,g,\qquad\operatorname{inj}(M,h)\geq L,\qquad|\operatorname{Ric}_{h}|_{h}\leq K.$
Here $\operatorname{inj}(M,g):=\inf_{p\in M}\operatorname{inj}_{g}(p)$ and $|\cdot|_{h}$ denotes the pointwise norm computed using $h$.
Lemma 1 (Metric comparison).
Let $g,h$ be Riemannian metrics on $M$ and let $C\geq 1$. If $C^{-1}g\leq h\leq Cg$, then for all $x,y\in M$, we have:
$C^{-1/2}\operatorname{dist}_{g}(x,y)\leq\operatorname{dist}_{h}(x,y)\leq C^{1/2}\operatorname{dist}_{g}(x,y),$
and pointwise on $M$,
$C^{-n/2}\,d\operatorname{vol}_{g}\leq d\operatorname{vol}_{h}\leq C^{n/2}\,d\operatorname{vol}_{g}.$
In particular, notice that if $(M,g)$ is complete, then $(M,h)$ is complete.
Proof.
For an arbitrary tangent vector $v$, it holds that:
$|v|_{h}^{2}=h(v,v)\leq Cg(v,v)=C|v|_{g}^{2},\qquad|v|_{h}^{2}\geq C^{-1}|v|_{g}^{2},$
so $|v|_{h}\leq C^{1/2}|v|_{g}$ and $|v|_{h}\geq C^{-1/2}|v|_{g}$. We now integrate along the curves and obtain the required distance bounds. For the volume forms, we choose a $g$-orthonormal basis so that $g$ is represented by $I$ and $h$ by a positive definite matrix $A$. It follows that $C^{-1}I\leq A\leq CI$, so we infer that $\det A\in[C^{-n},C^{n}]$ and $d\operatorname{vol}_{h}=\sqrt{\det A}\,d\operatorname{vol}_{g}$. We obtain completeness directly from the distance comparison. Observe that an $h$-Cauchy sequence is $g$-Cauchy, so it converges in $(M,g)$. In particular, it also converges in $(M,h)$ by the reverse comparison, which concludes the proof. ∎
Lemma 2 (Constant rescaling).
Let $\lambda>0$ and $\hat{g}:=\lambda g$. It follows that:
$\operatorname{inj}(M,\hat{g})=\sqrt{\lambda}\,\operatorname{inj}(M,g).$
Moreover, note that the Levi–Civita connections of $g$ and $\hat{g}$ are the same. In particular, we have:
$\operatorname{Ric}_{\hat{g}}=\operatorname{Ric}_{g}\quad\text{as $(0,2)$-tensors}.$
As a result, for an arbitrary $(0,2)$-tensor $T$, the following holds:
$|T|_{\hat{g}}=\lambda^{-1}|T|_{g}.$
Proof.
Lengths and distances scale by $\sqrt{\lambda}$ under $g\mapsto\lambda g$, and so does the injectivity radius. For constant rescaling, we have:
[…]
$=\frac{1}{2}(\lambda^{-1}g^{k\ell})(\lambda\partial_{i}g_{j\ell}+\lambda\partial_{j}g_{i\ell}-\lambda\partial_{\ell}g_{ij})$
$=\Gamma^{k}_{ij}(g),$
so the connections are the same. Notice that the $(1,3)$ curvature tensor is unchanged, and so is $\operatorname{Ric}$ as a $(0,2)$-tensor. Finally, we observe that $|T|_{\lambda g}^{2}=(\lambda g)^{ia}(\lambda g)^{jb}T_{ij}T_{ab}=\lambda^{-2}|T|_{g}^{2}$. ∎
Lemma 3 (Morrey–Sobolev on a unit ball).
Let $p>n$ and set $\alpha:=1-\frac{n}{p}\in(0,1)$. There exists a constant $C_{\mathrm{M}}=C_{\mathrm{M}}(n,p)>0$ such that for every $u\in W^{1,p}(B(0,1))$, there exists a representative (denoted by $u$) with $u\in C^{0,\alpha}(B(0,1))$ and
$u]_{C^{0,\alpha}(B(0,1))}\leq C_{\mathrm{M}}\|\nabla u\|_{L^{p}(B(0,1))}.$
Proof.
See, for instance, [5, p. 266]. ∎
Lemma 4 (Morrey estimate).
Let $p>n$ and set $\alpha=1-\frac{n}{p}$. Let $C_{\mathrm{M}}=C_{\mathrm{M}}(n,p)$ be the constant from Lemma 3. It follows that for every $r>0$ and every $u\in W^{1,p}(B(0,r))$,
$r^{\alpha}\,[u]_{C^{0,\alpha}(B(0,r))}\ \leq\ C_{\mathrm{M}}\,r^{1-\frac{n}{p}}\|\nabla u\|_{L^{p}(B(0,r))}.$
Proof.
The strategy is to scale to $B(0,1)$. Let us define $u_{r}(x):=u(rx)$. It follows that:
$\|\nabla u_{r}\|_{L^{p}(B(0,1))}$ $=r^{1-\frac{n}{p}}\|\nabla u\|_{L^{p}(B(0,r))},$
$[u_{r}]_{C^{0,\alpha}(B(0,1))}$ $=r^{\alpha}[u]_{C^{0,\alpha}(B(0,r))}.$
To conclude the proof, it suffices to apply Lemma 3 to $u_{r}$. ∎
Lemma 5.
Let $U\subset\mathbb{R}^{n}$ be open and let $m$ be a $C^{2}$ Riemannian metric on $U$ with coefficients $(m_{ij})$ in the standard coordinates. Let us assume that for some constants $c\geq 1$ and $A,B\geq 0$, it holds that:
$c^{-1}\delta\leq(m_{ij})\leq c\,\delta,\qquad\sup_{U}|\partial m_{ij}|\leq A,\qquad\sup_{U}|\partial^{2}m_{ij}|\leq B.$
We then know that there exists a constant $C_{\mathrm{curv}}=C_{\mathrm{curv}}(n,c)$ such that on $U$,
$|\operatorname{Rm}_{m}|_{m}\leq C_{\mathrm{curv}}\,(B+A^{2}).$
Proof.
Let $(m^{ij})$ be the inverse matrix. By ellipticity, we obtain $\sup_{U}|m^{ij}|\leq C_{0}(n,c)$. Furthermore, notice that the Christoffel symbols satisfy $|\Gamma|\leq C_{1}(n,c)A$. We now differentiate $m^{ip}m_{pj}=\delta^{i}_{j}$ and get $|\partial m^{ij}|\leq C_{2}(n,c)A$, so $|\partial\Gamma|\leq C_{3}(n,c)(A^{2}+B)$. The curvature components are precisely:
$R^{k}{}_{ij\ell}=\partial_{i}\Gamma^{k}_{j\ell}-\partial_{j}\Gamma^{k}_{i\ell}+\Gamma^{k}_{ip}\Gamma^{p}_{j\ell}-\Gamma^{k}_{jp}\Gamma^{p}_{i\ell},$
so $\sup|R^{k}{}_{ij\ell}|\leq C_{4}(n,c)(A^{2}+B)$. By ellipticity, we can then transform this to the invariant norm, as required. ∎
Lemma 6 (Croke).
There exists a constant $v_{n}>0$ that depends only on $n$ such that the following holds. If $(N^{n},m)$ is a compact Riemannian manifold without boundary, then for every $p\in N$ and every $0<r\leq\operatorname{inj}(N,m)/2$, it holds that:
$\operatorname{vol}_{m}\bigl(B_{m}(p,r)\bigr)\geq v_{n}\,r^{n}.$
Proof.
The result is due to [3]. ∎
Lemma 7 (Cheeger–Gromov–Taylor).
Let $(N^{n},m)$ be a connected complete Riemannian manifold with sectional curvature bounds $\lambda\leq\operatorname{sec}_{m}\leq\Lambda$. Let $p\in N$ and let $r>0$ with $r<\pi/(4\sqrt{\Lambda})$ if $\Lambda>0$. It then follows that:
$\operatorname{inj}_{m}(p)\ \geq\ r\,\frac{\operatorname{vol}_{m}(B_{m}(p,r))}{\operatorname{vol}_{m}(B_{m}(p,r))+V^{n}_{\lambda}(2r)},$
where $V^{n}_{\lambda}(\rho)$ denotes the volume of the radius-$\rho$ ball in the simply connected $n$-dimensional space form of constant sectional curvature $\lambda$.
Proof.
This is the first inequality in [4, Theorem 3.7]. ∎
We now have all the auxiliary lemmas in place. Let us now consider the proof of the main theorem.
Proof of Theorem 1.
We understand all inequalities between metrics in the sense of quadratic forms. For a $(0,2)$-tensor $T$, we compute the pointwise norm $|T|_{m}$ with respect to the metric $m$. If $M$ is not connected, we apply the argument below to each connected component. Given the fact that the hypotheses and the constants depend only on $(n,\ell,k)$, the constants we obtain are uniform. Therefore, we assume that $M$ is connected.
We first scale so that $\operatorname{inj}(M,\bar{g})\geq 1$. Note that compactness holds. Given that $\operatorname{Ric}_{g}\geq kg$ with $k>0$, we set H:=k/(n1)>0H:=k/(n-1)>0 so that $\operatorname{Ric}_{g}\geq(n-1)H\,g$. By Bonnet–Myers, we know that $(M,g)$ has finite diameter, and since it is complete, the Hopf–Rinow theorem implies that $(M,g)$ is compact (in equivalent terms, it is closed).
We set:
$\bar{g}:=\ell^{-2}g.$
By Lemma 2, we obtain:
$\operatorname{inj}(M,\bar{g})=\ell^{-1}\operatorname{inj}(M,g)\geq 1.$ (1)
It also holds that $\operatorname{Ric}_{\bar{g}}=\operatorname{Ric}_{g}$ as $(0,2)$-tensors, and so
$\operatorname{Ric}_{\bar{g}}\geq k\,g=k\ell^{2}\,\bar{g},$
so, in particular, $\operatorname{Ric}_{\bar{g}}\geq 0$ holds.
It suffices to prove the theorem for $(M,\bar{g})$ with $\operatorname{inj}(M,\bar{g})\geq 1$. Note that if $\bar{h}$ satisfies
$C^{-1}\bar{g}\leq\bar{h}\leq C\bar{g},\qquad\operatorname{inj}(M,\bar{h})\geq\bar{L},\qquad|\operatorname{Ric}_{\bar{h}}|_{\bar{h}}\leq\bar{K},$
then setting $h:=\ell^{2}\bar{h}$ and applying Lemma 2 gives
$C^{-1}g$ $\leq h\leq Cg,$
$\operatorname{inj}(M,h)$ $=\ell\,\operatorname{inj}(M,\bar{h})\geq\ell\bar{L},$
$|\operatorname{Ric}_{h}|_{h}$ $=\ell^{-2}|\operatorname{Ric}_{\bar{h}}|_{\bar{h}}\leq\ell^{-2}\bar{K}.$
We will now consider a conjugate-radius lower bound. To this end, for $p\in M$, let $\operatorname{conj\,rad}_{\bar{g}}(p)$ denote the conjugate radius at $p$. Observe that if $\exp_{p}$ is a diffeomorphism on the open $\bar{g}$-ball of radius $r$ in $T_{p}M$, then $\exp_{p}$ has no critical points there. This implies that $r\leq\operatorname{conj\,rad}_{\bar{g}}(p)$. We then obtain:
$\operatorname{inj}_{\bar{g}}(p)\leq\operatorname{conj\,rad}_{\bar{g}}(p)\qquad\forall p\in M.$
We apply (1) and get
$\operatorname{conj\,rad}_{\bar{g}}(p)\geq 1\qquad\forall p\in M.$ (2)
We now work towards establishing weak harmonic $C^{0,\alpha}$ control. Let us first fix
$p_{0}:=2n>n,\qquad\alpha:=1-\frac{n}{p_{0}}=\frac{1}{2}.$
Note that by (2) and $\operatorname{Ric}_{\bar{g}}\geq 0$, we may apply the results from [2] (if we take $H=0$ and $\rho=1$). More precisely, there exists a function $Q_{1}:(0,1]\to(0,\infty)$ with $\lim_{r\downarrow 0}Q_{1}(r)=0$ such that for every $r\in(0,1]$, the following is true:
$\|(M,\bar{g})\|_{W;hL^{1,p_{0}},r}\ \leq\ Q_{1}(r),$ (3)
where $\|\cdot\|_{W;hL^{1,p_{0}},r}$ denotes the weak harmonic $L^{1,p_{0}}$-norm on scale $r$ in the sense of [2]. We stress that this is the weak harmonic $L^{1,p_{0}}$-norm (not just the weak $L^{1,p_{0}}$-norm). The results from [2] provide the required harmonic coordinate charts. We also absorb absolute factors into $Q$. Finally, if we replace $Q_{1}$ by its monotone envelope, we may assume $Q_{1}$ is nondecreasing.
Let us fix $r\in(0,1]$. We now analyze the definition of $\|(M,\bar{g})\|_{W;hL^{1,p_{0}},r}$ [2] in more detail. (3) means that for each $p\in M$, there exists a local diffeomorphism $\varphi:B(0,r)\to U\subset M$ with $p\in U$ with the property that the pullback metric coefficients $(\bar{g}_{ij})=(\varphi^{*}\bar{g})_{ij}$ satisfy:
$e^{-Q_{1}(r)}\delta\leq(\bar{g}_{ij})\leq e^{Q_{1}(r)}\delta,\qquad r^{1-\frac{n}{p_{0}}}\|\partial\bar{g}_{ij}\|_{L^{p_{0}}(B(0,r))}\leq Q_{1}(r),$ (4)
in the sense of [2]. We keep the normalization $e^{\pm Q_{1}(r)}$.
In particular, the quadratic-form bounds imply that
$|\bar{g}_{ij}|\leq\sqrt{\bar{g}_{ii}\bar{g}_{jj}}\leq e^{Q_{1}(r)}\qquad\text{on }B(0,r).$
We now apply Lemma 4 componentwise to $\bar{g}_{ij}$ and obtain
$r^{\alpha}\,[\bar{g}_{ij}]_{C^{0,\alpha}(B(0,r))}\leq C_{\mathrm{M}}(n,p_{0})\,Q_{1}(r).$
It follows that
$r^{\alpha}\|\bar{g}_{ij}\|_{C^{0,\alpha}(B(0,r))}$ $\leq r^{\alpha}\|\bar{g}_{ij}\|_{L^{\infty}(B(0,r))}$
$\qquad+r^{\alpha}[\bar{g}_{ij}]_{C^{0,\alpha}(B(0,r))}$
$\leq r^{\alpha}e^{Q_{1}(r)}+C_{\mathrm{M}}Q_{1}(r).$
We now define, for $0<r\leq 1$,
$Q_{0}(r):=\max\Bigl\{\,Q_{1}(r),\ r^{\alpha}e^{Q_{1}(r)}+C_{\mathrm{M}}Q_{1}(r)\Bigr\}.$
We infer that $Q_{0}$ is nondecreasing and $\lim_{r\downarrow 0}Q_{0}(r)=0$. We extend $Q_{0}$ to $(0,\infty)$ by setting $Q_{0}(r):=Q_{0}(1)$ for $r\geq 1$. Note that this keeps $Q_{0}$ nondecreasing. Given that $Q_{0}(r)\geq Q_{1}(r)$ holds, we may apply the same harmonic charts as in (3). The charts are unchanged, so they remain harmonic. Moreover, the ball-containment condition in the definition of the weak norm is preserved, because if we replace $Q_{1}$ by $Q_{0}\geq Q_{1}$, we merely shrink the required ball. It follows that
$\|(M,\bar{g})\|_{W;hC^{0,\alpha},r}\ \leq\ Q_{0}(r)\qquad\forall r\in(0,1].$ (5)
So we have $(M,\bar{g})\in M(n,\alpha,Q_{0})$ in the sense of [2, Definition 1].
We now apply the smoothing theorem of Petersen–Wei–Ye [2, Theorem 1.1] to $(M,\bar{g})\in M(n,\alpha,Q_{0})$. To this end, fix $\varepsilon=1$ and $m=2$. We observe that Theorem 1.1 is uniform for all $0<r\leq 1$, and the parameter $r$ occurs only in the weak norms $\|\cdot\|_{W;C^{k,\alpha},r}$. We use the weak $C^{2,\alpha}$ bound at scale $r=1$, so we take $Q_{0}(1)$ in the constants. Theorem 1.1 is valid for an arbitrary integer $m\geq 2$. Most importantly, its smoothing construction [2] forms a smooth metric with uniform $C^{m,\alpha}$ bounds. In the argument below, we only use the $C^{2,\alpha}$ bounds (so we take $m=2$). We arrive at a Riemannian metric $\bar{h}$ on $M$ with the property that
$e^{-1}\bar{g}\leq\bar{h}\leq e\,\bar{g},$ (6)
and such that its weak $C^{2,\alpha}$-norm on scale $1$ is bounded:
$\|(M,\bar{h})\|_{W;C^{2,\alpha},1}\leq Q_{2},$ (7)
where $Q_{2}$ depends only on $n$, $\alpha$, $\varepsilon$ (here $1$), $m=2$, and $Q_{0}(1)$.
We now consider (7), that is, for each $p\in M$, there exist a local diffeomorphism $\varphi:B(0,1)\subset\mathbb{R}^{n}\to U\subset M$ and a point $\tilde{p}\in B(0,1)$ with $\varphi(\tilde{p})=p$ such that, if we write $(\bar{h}_{ij})=(\varphi^{\ast}\bar{h})_{ij}$ on $B(0,1)$, we get:
$e^{-Q_{2}}\delta$ $\leq(\bar{h}_{ij})\leq e^{Q_{2}}\delta,$ (8)
$\max_{|\beta|\leq 2}\sup_{B(0,1)}|\partial^{\beta}\bar{h}_{ij}|\;+\;[\partial^{2}\bar{h}_{ij}]_{C^{\alpha}(B(0,1))}$ $\leq Q_{2}.$ (9)
We next work towards determining the uniform curvature and Ricci bounds for $\bar{h}$. We fix $p\in M$ and choose $\varphi$ as in the smoothing construction above. We also apply Lemma 5 to $m=\varphi^{\ast}\bar{h}$ on $B(0,1)$, using (8)–(9). It follows that there exists $\Lambda=\Lambda(n,Q_{2})$ such that
$|\operatorname{Rm}_{\bar{h}}|_{\bar{h}}(p)\leq\Lambda.$
Given that $p$ was arbitrary, it follows that
$|\operatorname{Rm}_{\bar{h}}|_{\bar{h}}\leq\Lambda\qquad\text{on }M.$ (10)
Observe that since $\operatorname{Ric}$ is a contraction of $\operatorname{Rm}$, there exists $c_{\mathrm{Ric}}=c_{\mathrm{Ric}}(n)>0$ such that
$|\operatorname{Ric}_{\bar{h}}|_{\bar{h}}\leq c_{\mathrm{Ric}}\,|\operatorname{Rm}_{\bar{h}}|_{\bar{h}}\leq c_{\mathrm{Ric}}\,\Lambda\;=:\;\bar{K}.$ (11)
We also know that $|\sec_{\bar{h}}|\leq|\operatorname{Rm}_{\bar{h}}|_{\bar{h}}$, and so
$\Lambda\leq\operatorname{sec}_{\bar{h}}\leq\Lambda\qquad\text{on }M.$ (12)
We now obtain a uniform lower volume bound for $\bar{h}$-balls. We apply Lemma 6 to $(M,\bar{g})$ and use (1) to get, for all $p\in M$ and all $0<r\leq 1/2$,
$\operatorname{vol}_{\bar{g}}\bigl(B_{\bar{g}}(p,r)\bigr)\geq v_{n}\,r^{n}.$ (13)
We also apply the bi-Lipschitz comparison (6). By Lemma 1 with $C=e$, it follows that $d_{\bar{h}}\leq e^{1/2}d_{\bar{g}}$, so
$B_{\bar{g}}\bigl(p,e^{-1/2}r\bigr)\subset B_{\bar{h}}(p,r).$
We also have $d\operatorname{vol}_{\bar{h}}\geq e^{-n/2}d\operatorname{vol}_{\bar{g}}$. Therefore, for $0<r\leq 1/2$,
$\operatorname{vol}_{\bar{h}}\bigl(B_{\bar{h}}(p,r)\bigr)\geq v_{n}e^{-n}\,r^{n}.$ (14)
It remains to estimate the injectivity radius of $\bar{h}$ using Cheeger–Gromov–Taylor. We first apply Lemma 7 to $(M,\bar{h})$, use (12), and take $\lambda=-\Lambda$. Let us define:
$r_{*}:=\begin{cases}\min\bigl\{\,\frac{1}{2},\ \frac{\pi}{8\sqrt{\Lambda}}\,\bigr\},&\Lambda>0,\\[4.0pt] \frac{1}{2},&\Lambda=0.\end{cases}$
It follows that $0<r_{*}\leq 1/2$, and if $\Lambda>0$, we have $r_{*}<\pi/(4\sqrt{\Lambda})$.
By (14), we get:
$\operatorname{vol}_{\bar{h}}(B_{\bar{h}}(p,r_{*}))\geq v_{n}e^{-n}r_{*}^{n}.$
We infer that Lemma 7 gives, for every $p\in M$,
$\operatorname{inj}_{\bar{h}}(p)\ \geq\ r_{*}\,\frac{v_{n}e^{-n}r_{*}^{n}}{v_{n}e^{-n}r_{*}^{n}+V^{n}_{-\Lambda}(2r_{*})}\ =:\ \bar{L}.$
In particular, it is true that
$\operatorname{inj}(M,\bar{h})\geq\bar{L}>0.$ (15)
Finally, let us now return to the original scale. We set $h:=\ell^{2}\bar{h}$. By (6), we have
$e^{-1}g\leq h\leq e\,g,$
so we may take $C:=e$. By (15) and Lemma 2,
$\operatorname{inj}(M,h)=\ell\,\operatorname{inj}(M,\bar{h})\geq\ell\,\bar{L},$
so we take $L:=\ell\,\bar{L}$. Finally, by (11) and Lemma 2,
$|\operatorname{Ric}_{h}|_{h}=\ell^{-2}|\operatorname{Ric}_{\bar{h}}|_{\bar{h}}\leq\ell^{-2}\bar{K},$
so we set $K:=\ell^{-2}\bar{K}$. This concludes the proof. ∎
Remark 1.
Note that we used the Bonnet–Myers theorem only to obtain compactness, which, in turn, was required to apply Croke’s result. In the rest of the proof, we used $\operatorname{Ric}_{\bar{g}}\geq 0$ with the lower injectivity radius bound. If we were to assume compactness separately and still assume $\operatorname{Ric}_{\bar{g}}\geq 0$ and $\operatorname{inj}_{\bar{g}}\geq 1$ (in other terms, $\operatorname{Ric}_{g}\geq 0$ and $\operatorname{inj}_{g}\geq\ell$ before scaling), then we could take the constants to depend only on $n$ and $\ell$.
Statements and Declarations
Competing Interests
The author declares that there are no competing interests.